OpenStudy (twistrose):

What is the value of x? (6+8x)/2=5x –2–2 –1–1 1313 3

OpenStudy (rcp031):

(6+8x)/2=5x 3+4x=5x x=3

Awolflover1 (awolflover1):

let's go step by step with the user please.

Awolflover1 (awolflover1):

\[\frac{ 6+8x }{ 2 }=5x\] right ?

OpenStudy (nevillesnetwork):

Step 1: Simplify both sides of the equation. 6+8x 2 =5x Simplify: 4x+3=5x Step 2: Subtract 5x from both sides. 4x+3−5x=5x−5x −x+3=0 Step 3: Subtract 3 from both sides. −x+3−3=0−3 −x=−3 Step 4: Divide both sides by -1. −x −1 = −3 −1 x=3 Answer: x=3

OpenStudy (nevillesnetwork):

Welcome! Medal and fan plz :3 ty!

Awolflover1 (awolflover1):

copy and paste i see :|

OpenStudy (twistrose):

thanks guys

OpenStudy (nevillesnetwork):

I cant copy and paste?

OpenStudy (nevillesnetwork):

I dont think thats a problem....